OpenStudy (cwtan):

How to use Row echelon form of matrix solving -2x+y=-3 x-4y=-2

OpenStudy (anonymous):

Easy way to solve system of equations you get the value of one variable from the last row and work in reverse to find the values of the other unknown variables.

OpenStudy (cwtan):

How? For example -2x+y=-3 and x-4y=-2 Can explain for me step by step?

OpenStudy (anonymous):

You have a system just :)

OpenStudy (anonymous):

-2x+y=-3 x-4y=-2 _________________-- -2x+y=-3 2(x-4y)=2(-2) ___________________- -2x+y=-3 2x-8y=-4 -7y=-7 y=1 x=2 Just this :)

OpenStudy (anonymous):

@sami-21 : Right ?:)

OpenStudy (anonymous):

yes correct .

OpenStudy (anonymous):

but its not echelon form.

OpenStudy (anonymous):

there are some fundamental row operations 1) you can interchange rows 2) you can multiply or divide the row with real no 3) you can add subtract multiple of any row in another row . writing the equations in matrix form

OpenStudy (anonymous):

yeah and I do *2 to all :)

OpenStudy (cwtan):

Oh Thanks~

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

there are some fundamental row operations 1) you can interchange rows 2) you can multiply or divide the row with real no 3) you can add subtract multiple of any row in another row . writing the equations in matrix form